50, 000 J of heat was absorbed by the surroundings when 150 kJ of

work was done by the system.


1
Expert's answer
2021-06-28T08:11:19-0400

ΔU=q+w=150−50=100kJ


Therefore, internal energy change increases by 100 kJ.